# Understanding long-term energy use in off-grid solar home systems in sub-Saharan Africa

### v1

@@ V-B2 Cluster space trend
TABLE III: Mean characteristics of households, separated by their dominant cluster – the cluster in which they spend the majority of their time.
Figure 4: All the households are split into their dominant cluster groups (y-axis). For each of these groups, the spread of days in the other clusters is found (x-axis).
Figure 5: Proportion of households in each cluster, across the lifetime of all the systems (top). The bottom subplots split the households into high energy ($>50$ Wh of appliances) and low energy ($\leq 50$ Wh of appliances). The black dashed line indicates when the number of households in the sample goes below 10.
@@
The higher consumption clusters (single peak and double peak) have lower homogeneity. Table III shows that households with higher consumption dominant clusters, have greater appliance power, with the vast majority of households owning a TV. Therefore, households with higher consumption dominant clusters have the potential to be in all clusters, however, the majority of households with other dominant clusters do not have the potential to consume the energy in the higher consumption clusters due to lack of appliances. Additionally, in the higher consumption clusters, households have more flexible appliances. The base households have lights and phone chargers – these are relatively inflexible loads, as households need lighting when it is dark and to charge phones when the batteries have run out. For this work, we define anything other than lighting and phone charging to be a flexible load. In the higher consumption clusters, households have appliances like TVs and radios which they may not want to use every day, leading to changes in the household energy consumption pattern more often.
@@
Fig. 4 depicts how households, split into their dominant clusters, consume energy when not following their dominant cluster’s behaviour. This further demonstrates the possibility of higher consumption clusters to be in all cluster spaces, whereas is it very uncommon for moderate use households to be in single peak and double peak.
@@
Economic outage is observed in all clusters, but particularly prevalent in the three lower consumption clusters (low use, moderate use and nighttime use). This is logical, as customers who initially pay less for the system have fewer appliances and can be suspected to be more likely to face financial hardship, or be less engaged with the system – both of which lead to missed payments and economic outages.
@@
Beyond the common affinity of all clusters to the economic outage cluster, Fig. 4 shows that clusters with similar magnitude of consumption are more closely linked within households. Households with dominant clusters moderate use and nighttime use are most likely to be in each other’s cluster. Similarly, households with dominant cluster single peak and double peak are most likely to be in each other’s cluster. Therefore, while households energy consumption is clearly non-homogenous, it is more common for households to vary their consumption behaviour within similar magnitudes of use or to not use the system at all and not pay.
@@
Fig. 5 shows the proportion of households in each cluster on each day for the lifetime of the systems. The most important finding is the sharp increase in economic outage over the first year. This is the primary explanation for the reducing electricity consumption, as previously seen in Fig. 3, and shows that customers are reducing engagement with the SHS for economic reasons – either they cannot afford to keep up with payments or they are choosing to not spend their money on SHS credit. Once economic outage begins to level, low use starts to gradually increase. The low use cluster is important for showing when households are choosing to hardly use the system regardless of economic considerations. The increase therefore shows disengagement with the system for non-economic reasons.
@@
Previously, from Fig. 4, the link between nighttime use and moderate use was observed. This link is reinforced in Fig. 5 in the low energy household subplot, where nighttime use significantly reduces over the time of analysis, while moderate use increases. We hypothesise that this is the driver for the decreasing use in low energy households, found in Fig. 3 even when economic outages are ignored. Given their already low consumption levels, the shift from nighttime use to moderate use has a substantial impact. For instance, households might initially use lighting throughout the night for security purposes, but over time, reduce this to only evening hours. To fully understand the reasons behind this behavioural change, further research, including surveys, is necessary.
@@
The high energy household subplot in Fig. 5 shows a similar trend of consistently decreasing single peak and double peak with increase in moderate use and low use. Thus, households are reducing their peak consumption overtime and reducing the use of their flexible loads.
@@ V-C Utilisation rate and electricity use
Figure 6: Proportion of households in each cluster over time, split into households with a utilisation rate over 90% and under 90% at time of analysis. The black dashed line indicates when the number of households in the sample goes below 10. The bar chart shows the proportion of days in each cluster for households, grouped by their utilisation rate.
Fig. 5 illustrates a rise in economic outage over time. However, a deeper understanding of household behaviour can be gained by examining utilisation rates. Fig. 6 presents the temporal variation in the proportion of households within each cluster, segmented by utilisation rate. 530 of the households in the dataset have a utilisation rate of 0.9 and above, and only 121 households have a utilisation rate of less than 0.5. Since utilisation rates fluctuate over time, the rate assigned to each household reflects its level at the time of data collection.
The observed increase in economic outage is primarily driven by households with lower utilisation rates, which, by definition, is less evident in households with higher utilisation rates. Beyond economic outages, the cluster space looks different for different utilisation rates, and therefore the utilisation rate influences cluster membership.
In the low utilisation rate group, nighttime use significantly reduces, particularly within the first year. Low use and moderate use fluctuate but stay relatively constant, whereas single peak and double peak steadily decrease. There isn’t a significant increase in any clusters aside from economic outage in the lower utilisation households, so households here are clearly changing their energy behaviour due to financial reasons – either an affordability issue or a lack of desire to spend their money on the system.
For households with higher utilisation rate, the nighttime use continually decreases but at a steadier rate than in lower utilisation households. Higher utilisation rate households are more likely to be in single peak and double peak than lower utilisation rate households. The mean proportion of households in these two highest consumption clusters across the whole timeframe is 19% for low utilisation rate households and 32% for high utilisation households. Thus, households that pay more consistently, are more likely to be in the higher use clusters. However, beyond the first year, single peak and double peak rates decrease, while low use and moderate use increase. The increase in these lower use clusters demonstrates how households are reducing their electricity consumption when economic outage is negligible, which further builds on the previously stated hypothesis of households decreasing electricity consumption while still paying for and engaging with the SHS.
Overall, splitting the households by their utilisation rates draws several key conclusions.
@@
  • Economic outage increases significantly over the first year of ownership in low utilisation households. In high utilisation households, the small proportion of economic outages increases gradually over time, peaking at the 3-year mark.
  • There is a shift to lower consumption clusters over time, which is more pronounced in households with utilisation rate over 90%.
@@
VI Conclusion
This work has analysed a large dataset of SHS across sub-Saharan Africa. The DTW-MIP clustering effectively found 5 distinct daily load profiles, archetypal defines as low use, moderate use, nighttime use, single peak and double peak. The representation of these across the dataset shows the SHS is only fully utilised, as in single peak and double peak, on 24% of days – the same proportion of days that the SHS are in economic outage and low use. The majority of time, the SHS provide the basic provisions of moderate use and nighttime use. Households frequently change how they use their SHS, with cluster homogeneity for households averaging 55%. This highlights the importance of granular load demand data, instead of using a one-size-fits-all approach.
Analysis of household behaviour over a period of up to 1,200 days revealed a general decrease in energy consumption among both high and low consumption households. While economic outages are not uncommon under the PAYG scheme, they are not the sole reason for decreasing energy consumption, indicating that changes in energy behaviour extend beyond financial considerations. The decrease in energy consumption when not in economic outage is characterised by a decrease of nighttime energy use, and reduction of days with large evening peaks. Utilisation rates are slightly higher for households that predominantly exhibit single peak and double peak profiles, but even these households—characterised by more flexible appliances and TV ownership, meaning more expensive daily rates—experience economic outages.
Overall, the systems are often underused and therefore oversized, leading to knock-on effects on costs for both households and suppliers. The significant proportion of economic outages suggests that costs remain a limitation for many households, warranting further investigation into how maximising energy provision from each SHS could help reduce the overall cost of energy.
Further work should include surveys of households to understand the underlying causes of the decrease in energy consumption while the systems are still being paid for. Moreover, continuous study of energy behaviour as households gain access to electricity is essential for better understanding their needs and how these evolve over time.
@@ VII Acknowledgements
The authors thank BBOXX for project funding and access to SHS data, and EPSRC for project funding. Prof. Parikh is grateful to the Royal Academy of Engingeering for part funding.

### v2

@@ V-B2 Cluster allocation trend
TABLE III: Household mean characteristics separated by dominant cluster—where they spend the majority of time.
@@
Higher consumption clusters (single peak and double peak) exhibit lower homogeneity. Table III also shows that households with days in higher consumption clusters have greater appliance power, with the majority owning a TV. Therefore, households with days in higher consumption dominant clusters have the potential to be in all clusters, however, the majority of households with other dominant clusters do not consume larger amounts of energy due to lack of appliances. Additionally, in the higher consumption clusters, households have more flexible loads. Base households have lights and phone chargers and these are relatively inflexible because lighting is needed when it is dark and phone charging when batteries are depleted. Here, we define anything other than lighting and phone charging to be a flexible load. In the higher consumption clusters, households have appliances like TVs and radios which they may not use every day, leading to more frequent day-to-day changes in electricity consumption.
@@
Fig. 4 depicts how households, split into their dominant clusters, consume energy when not following the dominant cluster behaviour. This further demonstrates that households dominated by higher consumption can also have usage days in all other clusters, whereas is it very uncommon for moderate use dominated households to have days in the single peak and double peak clusters.
Figure 4: Spread of usage over clusters: for each household with a certain dominant cluster (rows), the proportion of time spent in other clusters is indicated by the columns.
@@
Economic outage is observed in all households, but is particularly prevalent in the three lower consumption dominated cases (low use, moderate use and nighttime use). This is logical, because customers who initially pay less have access to smaller amounts of available energy, and fewer appliances, and may be more likely to face financial hardship or be less engaged with their system—both of which lead to missed payments and economic outages.
@@
Fig. 4 also shows that clusters with similar average consumption are more closely linked within households (e.g., moderate use and nighttime use, similarly single peak and double peak). Therefore, while household energy consumption is clearly non-uniform, it is more common for households to vary their consumption behaviour within similar magnitudes of use or to not use the system at all and not pay.
@@
Fig. 5 shows the proportion of households in each cluster on each day. The most important finding is the sharp increase in economic outage over the first year. This is the primary explanation reduced electricity consumption, as previously seen in Fig. 3, and shows that customers either cannot afford to keep up with payments or are choosing not to spend their money on SHS credit. Once economic outage begins to plateau, low use starts to gradually increase. The low use cluster shows when households are choosing to use the system minimally regardless of economic considerations. The increase therefore shows disengagement with the system for non-economic reasons.
Figure 5: Top: Proportion of households in each cluster, every day. Bottom: High-use households ($>50$ Wh of appliances), left, and low-use ($\leq 50$ Wh of appliances), right. Black dashed line indicates when number of households drops below 10.
@@
Previously, Fig. 4, a link between nighttime use and moderate use was observed. This is reinforced in Fig. 5, especially in the low-use subplot, where nighttime use significantly reduces over the time of analysis, while moderate use increases. We hypothesise that this is the driver for the decreasing use in low-use households shown in Fig. 3. Given their already low consumption, a shift from nighttime use to moderate use has a substantial impact. For instance, households might initially use lighting throughout the night for security purposes, but over time, reduce this to evening hours only. To fully understand the reasons behind this behavioural change, further research including surveys is necessary.
@@
The high energy household subplot in Fig. 5 shows a similar trend of consistently decreasing single peak and double peak with increase in moderate use and low use. Thus, households reduce peak consumption over time and reduce the use of flexible loads.
@@ V-C Utilisation rate and electricity use
Figure 6: Proportion of households in each cluster over time, split into households with a utilisation rate over 90% and under 90% at time of analysis. The black dashed line indicates when the number of households in the sample goes below 10.
Fig. 5 illustrates a rise in economic outage over time. However, a deeper understanding of household behaviour can be gained by examining utilisation rates. Fig. 6 presents the time variation in the proportion of households within each cluster segmented by utilisation rate. There are 530 households in the dataset with a utilisation rate of 0.9 and above, but only 121 households with a utilisation rate of less than 0.5.
The observed increase in economic outage is primarily driven by households with lower utilisation rates, which, by definition, is less evident in households with higher utilisation rates. In the low utilisation rate group, nighttime use significantly reduces, particularly in the first year; low use and moderate use fluctuate but stay relatively constant, whereas single peak and double peak steadily decrease. In low utilisation households, there isn’t a significant increase in any cluster aside from economic outage, so these households are clearly changing their energy behaviour for financial reasons—either an affordability issue or a lack of desire to spend money on the system.
For households with higher utilisation rate, the proportion of nighttime use continually decreases but at a steadier rate than in lower utilisation households. Higher utilisation rate households are more likely to be in single peak and double peak clusters than lower utilisation rate households. The mean proportion of households in these two high-consumption clusters across the whole time period is 19% for low utilisation rate households and 32% for high utilisation households. Thus, households that pay more consistently are more likely to be in the higher use clusters. However, beyond the first year, single peak and double peak rates decrease, while low use and moderate use increase. This demonstrates how households are reducing their electricity consumption even when economic outage is negligible, which further builds on the previously stated hypothesis that households decrease electricity consumption even while still paying for and engaging with the SHS.
Overall, splitting the households by utilisation rates highlights several key issues:
@@
  • Economic outage increases significantly over the first year of ownership in low utilisation households. In high utilisation households, the small proportion of outages increases gradually over time, peaking at the 3-year mark.
  • There is a shift to lower consumption clusters over time and it is more pronounced in households with utilisation rate over 90%.
@@
VI Conclusions
This work has analysed a large dataset of energy use in 1,000 solar home systems across sub-Saharan Africa. Time series clustering identified five distinct daily usage profiles, archetypally defined as low use, moderate use, nighttime use, single peak and double peak, in order of increasing mean energy use. The representation of these across the dataset shows that systems are only fully utilised, i.e., their energy usage is in the single peak or double peak cluster, on 24% of days—the same proportion of days that the systems are in the economic outage and low use clusters. The majority of time, the systems provide basic supply of moderate use and nighttime use electricity. Households frequently change, from day to day, how they use their systems, with cluster homogeneity for households averaging only 55%.
Analysis of trends in household behaviour over a period of up to 1,200 days revealed a general decrease in energy consumption among both high and low consumption households. While economic outages are not uncommon under the PAYG scheme, they are not the sole reason for decreasing energy consumption, indicating that changes in energy behaviour extend beyond financial considerations. The decrease in energy consumption when not in the economic outage state is characterised by a decrease in nighttime energy use, and reduction of days with large evening peaks. Utilisation rates are slightly higher for households that predominantly exhibit single peak and double peak profiles, but even these households experience economic outages.
Overall, the systems are often underused and therefore oversized, leading to knock-on effects on costs for households and suppliers. The significant proportion of economic outages suggests that costs remain a limitation for many households, warranting further investigation into how to maximise energy provision from each SHS in order to reduce the overall cost of energy. There is a need for SHS designs and pricing structures to better align with the actual energy requirements of households. Adaptable solutions may be key for long term adoption, and this may include downsizing systems over time in some cases.
Further work should include surveys of households to understand the underlying causes of the non-financial decrease in energy consumption. Moreover, continuous study of energy behaviour as households gain access to electricity is essential for better understanding their needs and how these evolve over time.
@@ VII Acknowledgements
The authors thank BBOXX for project funding and access to SHS data, and EPSRC for project funding (ref. EP/W027321/1). Prof. Parikh is grateful to the Royal Academy of Engineering for part funding.
For the purpose of Open Access, the authors apply a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.